OpenStudy (anonymous):

integrate 4x-1/(x^2 +2)(x-1)

OpenStudy (phi):

use partial fractions ?

OpenStudy (anonymous):

Partial fractions will work. \[\frac{4x-1}{(x^2+2)(x-1)}=\frac{Ax+B}{x^2+2}+\frac{B}{x-1}\]

OpenStudy (anonymous):

*That last B should be a C